OpenStudy (anonymous):

how would i simplify 2 to the third power+(8-5)*4-5to the second power ?

OpenStudy (anonymous):

do u mean this : 2^(3 + (3)^4 - 5) ^2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

2^3 + ((8-5)*4 -5)^2 is what is asked for if I understand it. So 8 + (12-5)^2 = 8 + 7^2 = 57.